# SAM-LAD: Segment Anything Model Meets Zero-Shot Logic Anomaly Detection

### v1

@@ II-A Semantic Segmentation
The field of semantic segmentation has witnessed significant advancements, with novel methods being developed that target improved precision and versatility. The Segment Anything Model (SAM)[16], a newly introduced framework, is revolutionizing object segmentation by leveraging both large-scale training datasets and innovative model architectures. Existing studies have explored various techniques, such as transformers with self-attention mechanisms[19], as well as methods that integrate diverse image contexts to enhance segmentation performance. Prior to SAM, notable segmentation models like Mask R-CNN[20] and DeepLab[21] have paved the way in extending segmentation capabilities through region proposals, feature pyramid networks, and atrous convolutions. SAM continues this trend of innovation by adopting a prompt-driven architecture that not only enhances traditional segmentation tasks but also explores uncharted territories in segmentation capabilities. Furthermore, interactive segmentation methods have previously been explored to address the ambiguity of object boundaries by enabling user inputs to guide segmentation predictions. However, SAM’s approach differs by providing a flexible way to specify segmentation prompts (e.g., points, boxes, masks) and generalize the segmentation to diverse scenarios.
In this work, we use SAM’s robust segmentation capabilities and generalization performance. By setting segmentation thresholds, we can effectively isolate the key objects to be detected from the scene. Consequently, our framework achieves a groundbreaking analysis of the entire scene, all based on zero-shot learning.
@@ II-B Keypoint Matching
Keypoint matching is a fundamental task in computer vision, crucial for applications such as object recognition, 3D reconstruction, and image stitching. The long-standing interest in this problem has resulted in several approaches over the years. Early classical methods like SIFT[22] and ORB[23] laid the groundwork by detecting handcrafted keypoint descriptors. These algorithms focused on creating robust descriptors that were invariant to scale, rotation, and partial occlusion, which made them effective for a wide range of tasks, despite being sensitive to image illumination changes. In recent years, deep learning techniques have revolutionized keypoint matching by learning discriminative feature representations from large datasets. CNNs have been widely adopted in keypoint detection and matching, such as the SuperPoint model[24], which directly learns keypoint detection and descriptor generation in an end-to-end fashion. Deep learning-based approaches offer much-improved robustness and generalization compared to handcrafted methods. More recent advances employ attention mechanisms to improve keypoint matching under challenging environments. Networks like SuperGlue[25] harness graph neural networks along with self- and cross-attention mechanisms to establish robust keypoint correspondences. The introduction of transformers into keypoint matching has shown further potential in improving the model’s context understanding, enabling them to handle complex matching scenarios more proficiently.
Inspired by SuperPoint and Superglue, we consider each object in the scene as a keypoint, and compress the keypoint features into a single feature vector using the proposed dynamic channel graph attention (DCGA) mechanism. Subsequently, we compute the object-to-object matching relationship using the concept of optimal transport.
@@ II-C Unsupervised Anomaly Detection
The core part of unsupervised anomaly detection algorithms lies in extracting prior knowledge of anomaly-free objects, serving as a criterion for anomaly detection tasks. Typically, current unsupervised anomaly detection approaches can be classified into two categories based on the type of anomalies they address: methods targeting structural anomalies, and methods targeting logical anomalies. Presently, the majority of anomaly detection methods are designed for structural anomalies, primarily using a robust feature extraction network to obtain high-level semantic features of a test image. The distance between features of test images and anomaly-free images is calculated, and the areas with large distances are determined to be anomalies. The existing methods usually use deep CNN models, such as ResNet[26] and EfficientNet[27], for feature extraction. For instance, Roth et al.[10] proposed Patchcore to detect anomalies on objects with a concise background. ResNet-50[26] pretrained on ImageNet[28] was selected as the feature extractor. However, these methods falter when detecting logical anomalies, as the local structures within the image appear normal. Hence, an increasing number of researchers are directing their focus toward logical anomaly detection. For instance, Zhang et al. have proposed DSKD[15], and Batzner has introduced EfficientAD[29]. Both methods are based on the Student-Teacher network and anomalies are obtained by comparing teacher and student differences. Leveraging a local-global branching approach, Yao et al. presented GLCF[13], while Bergmann et al. proposed GCAD[11] in a bid to understand the logical semantic constraints governing the entire image scene. Guo et al. presented THFR[14], which is based on template-guided reconstruction. These methods have exhibited promising efficacy in detecting certain logical anomalies, such as missing components and unexpected excess. However, their detection performance significantly drops when confronted with more complex logical anomalies, such as misordering, mismatches, and haphazard arrangement of objects.
The proposed SAM-LAD framework builds a zero-shot method without strenuous efforts to comprehend the logical semantics of the entire scene. By introducing SAM and proposing object matching model and anomaly measurement model, a global logical anomaly detection system was realized, perfectly addressing the limitations encountered by the existing methods mentioned above.
@@ III-E1 Dynamic Channel Graph Attention
This graph structure consists of two parts. The first part targets each object’s feature vector $(1,C)$. Specifically, two independent $C\times C$ matrices, that is, $A_{0}$ and $A_{1}$, constitute the adjacency matrices, representing the dependency relationships among channel vertices. $A_{0}$ is generated by a 1-D convolution layer and a Softmax function. It is a self-attention-based diagonal matrix designed to suppress irrelevant features. $A_{1}$ is a predefined identity matrix, representing only the vertex itself, and requires normalization. Thus, the adjacency matrix can be represented as:
$A=A_{0}\times A_{1}$ (4)
@@
$A_{2}$ $=\begin{bmatrix}0&S_{12}&\cdots&S_{1j}\\ S_{21}&0&\cdots&S_{2j}\\ \vdots&\vdots&\ddots&\vdots\\ S_{i1}&S_{i2}&\cdots\ &S_{ij}\\ \end{bmatrix},$ (5)
$S_{ij}$ $=\frac{S_{ij}^{\prime}}{\sum_{k=1}^{M}S_{ik}},$
$S_{ij}^{\prime}$ $=\hat{v}_{i}\cdot\hat{v}_{j}=\sum_{d=1}^{C}\hat{v}_{id}\cdot\hat{v}_{jd}.$
Where $S_{ij}^{\prime}$ represents the similarity between two object vertex in the M and $S_{ij}$ is its normalization. Note that, when generating the adjacency matrix, the weights updating is based solely on the similarity between one object vector and others, self-similarity is set to zero, that is, the diagonal of $A_{2}$ is zero.
@@ III-F Anomaly Measurement Module
Specifically, based on the indices from the $\mathbf{P}$ and trash bins, we calculate the difference of feature distribution for the object feature maps that were successfully matched with the corresponding object feature maps in the $k$ reference images to create the matching score map. We also calculate the difference of feature distribution for the unmatched object feature maps with the most closely matched object feature maps in the $k$ reference images to create the non-matching score map. The final score map was then obtained by adding these two score maps together. The advantage of this approach is that after matching at the object level, subsequent pixel-level computations allow not only the detection of logical anomalies but also a more detailed check for structural anomalies within each class of object.
@@ IV-B Evaluation metrics
IV-C Implementation Details
IV-D Comparison With the State-of-the-Art Models
IV-D1 Comprehensive Comparison
@@ IV-D2 MVTec LOCO AD
The comparative results are presented in Table II. Compared to the best baseline method and the SoTA method, our framework improves +25.6 and +0.7 on image-level AUROC, +45.0 and +5.4 on pixel-level sPRO, respectively. It is worth mentioning that, among all the comparison methods, only our approach implemented zero-shot while simultaneously maintaining the best detection performance. Furthermore, Fig. 8 displays several detection results of SAM-LAD on both structural and logical anomalies within the MVTec LOCO AD dataset, illustrating that our framework is adept at precisely pinpointing both types of anomalies.
IV-D3 MVTec AD
Beyond affirming the performance of our SAM-LAD in detecting logical anomalies, we have further appraise its capability to identify structural anomalies on the MVTec AD dataset, which exclusively encompasses structural anomalies. To be specific, given that each subset within the MVTec AD dataset features a scene with a single object against a regular background, we facilitate the segmentation into this single object by employing the SAM filter for subsequent matching and computation. Owing to the one-to-one fixed matching paradigm, the ultimate detection results are directly yielded by the AMM, thereby serving as an indirect assessment of the efficacy of our proposed AMM.
IV-D4 DigitAnatomy
IV-E Further Analysis
The outstanding logical anomaly detection results of our framework are based on the good performance of the Object Matching Model (OMM). Therefore, to verify the reasons behind the exceptional performance, for each category in MVTec LOCO AD and DigitAnatomy dataset, we compile a confusion matrix of the object matching results in Table V and Table VI, respectively. The ground truth is categorized by the number of objects in each test image that should be correctly matched and those that should not be matched. We use precision, recall, and F1 score as the metrics to evaluate the OMM’s performance.
The results in Table V and Table VI demonstrate the strong matching capability of the OMM, validating its effectiveness. However, in Table V, the matching accuracy for the pushpins and splicing connectors categories drops significantly. This is because, in these scenarios, each object is very similar, posing a considerable challenge for the object matching model. However, we did not directly use the unmatched object masks as the anomaly score. Instead, we design the AMM (Anomaly Measurement Module) to further calculate anomalies in the objects. This novel design greatly reduces the proposed SAM-LAD framework’s dependency on matching accuracy, as demonstrated by the final anomaly detection results in Section IV-D. A selection of the representative samples from the visualized matching results is shown in Fig. 13.
@@ IV-F1 Impact of the Backbone
We use the pre-trained backbone as the feature extractor in the proposed SAM-LAD. In this study, we primarily considered three different pre-trained backbone networks: the Wide-ResNet50 of the CNN type, the Swin-transformer of the ViT type, and the DINOv2 of the ViT type. The main results are presented in Table VII, where it is distinctly that DINOv2 manifests the most superior performance. Therefore, we adopt DINOv2 as the feature extractor for the proposed framework.
IV-F2 Impact of the FeatUp Configuration
Upon obtaining the feature maps from the images, we implement the FeatUp operation to upsample these maps and thus restore the spatial information that had been lost. In the study of FeatUp[18], the authors proffered four configurations of upsampling: 2$\times$, 4$\times$, 8$\times$, 16$\times$, and 32$\times$, which are displayed in Fig. 14. To verify the efficacy of FeatUp and to assess the impact of different upsampling configurations on the performance of anomaly detection, we conduct an ablation study focusing on the upsampling factors.
Figure 13: The matching results for each subset in MVTec LOCO AD and DigitAnatomy dataset, where the green lines indicate matching objects.
The main results are illustrated in the Table VIII. In comparison with the original feature maps, FeatUp significantly enhances detection capabilities. More specifically, as the factor of upsampling escalates, the detection and segmentation performance of SAM-LAD are concurrently fortified. Nevertheless, when the feature maps are subjected to upsampling by 16$\times$ and 32$\times$, there is dramatic inflation in the model’s FLOPs and a precipitous deterioration in inference speed, circumstances that are indefensible in pragmatic industrial settings. Consequently, to harmonize detection performance with the imperative of real-time inference, an 8$\times$ upsampler was opted for in our framework to restore the spatial information lost in the feature maps for use in subsequent computational processes.
IV-F3 Impact of the DCGA
To validate the effectiveness of the proposed DCGA, an ablation experiment was conducted without the incorporation of DCGA. The inception of DCGA was intended to extract high-dimensional features into a single feature vector, thereby simplifying the computation of similarity between objects. Pooling represents the most extensively utilized method for feature extraction. Consequently, we compare DCGA with both global average pooling (GAP) and global max pooling (GMP), with results delineated in Table IX. It can be observed that GMP outperforms GAP because only certain regions of the object feature map possess discernible feature values, while the background is predominantly nullified. Therefore, in such scenarios, GMP is more effective in capturing object features. Nevertheless, simply selecting the maximum value for pooling is insufficient. The introduction of DCGA has greatly enriched the description of object features. Utilizing the principles of graph neural networks, DCGA not only extracts efficacious object features from the high-dimensional channels but also captures the interrelationship information of features among objects within the same scene. This substantially improves the matching capabilities of the optimal-matching model and the ultimate performance of anomaly detection, thus confirming the validity of the proposed DCGA.
TABLE V: The Precision, Recall, and F1 Score of the Object Matching Module in the MVTec LOCO AD dataset.
breakfast box
Actual Predicted Positive Negative Precision: 99.2%
Positive 1502 19 Recall: 98.8%
Negative 12 117 F1 Score: 99.0%
juice bottle
Actual Predicted Positive Negative Precision: 99.5%
Positive 1503 26 Recall: 98.3%
Negative 8 110 F1 Score: 98.9%
pushpins
Actual Predicted Positive Negative Precision: 89.0%
Positive 3336 505 Recall: 86.9%
Negative 413 510 F1 Score: 87.9%
screw bag
Actual Predicted Positive Negative Precision: 99.5%
Positive 2426 28 Recall: 98.9%
Negative 13 261 F1 Score: 99.2%
splicing connectors
Actual Predicted Positive Negative Precision: 91.1%
Positive 2471 339 Recall: 87.9%
Negative 242 508 F1 Score: 89.5%
Mean Precision 95.66%
Mean Recall 94.16%
Mean F1 Score 96.52%
TABLE VI: The Precision, Recall, and F1 Score of the Object Matching Module in the DigitAnatomy.
Actual Predicted Positive Negative Precision: 96.5%
Positive 2853 147 Recall: 95.1%
Negative 102 1398 F1 Score: 95.8%
TABLE VII: Ablation Experiments of the Different Backbone on the MVTec LOCO AD Dataset.
Metrics Network Wide-ResNet50 Swin-Transformer DINOv2
@@ IV-F3 Impact of the DCGA
TABLE VIII: Ablation experiments of the Different FeatUp’s Upsampling Factors on the MVTec LOCO AD Dataset.
@@
TABLE IX: Ablation Experiments of the Different Feature Compress Methods on the MVTec LOCO AD Dataset.

### v2

@@ II-A Semantic Segmentation
Semantic segmentation has seen significant advancements, with novel methods targeting improved precision and versatility. The Segment Anything Model (SAM)[16] revolutionizes object segmentation by leveraging large training datasets and innovative architectures. Existing studies have explored techniques like transformers with self-attention[19] and methods that integrate diverse image contexts to enhance segmentation. Before SAM, models like Mask R-CNN[20] and DeepLab[21] extended segmentation capabilities using region proposals, feature pyramid networks, and atrous convolutions. SAM continues this innovation with a prompt-driven architecture that enhances traditional segmentation tasks and explores new segmentation capabilities. Interactive segmentation methods have addressed object boundary ambiguity by enabling user inputs to guide predictions. However, SAM differs by providing flexible segmentation prompts (e.g., points, boxes, masks) and generalizing segmentation to diverse scenarios.
In this work, we leverage SAM’s robust segmentation capabilities and generalization performance. By setting segmentation thresholds, we effectively isolate objects from the scene. Thus, our framework achieves groundbreaking scene analysis based on zero-shot learning.
@@ II-B Keypoint Matching
Keypoint matching is essential in computer vision for object recognition, 3D reconstruction, and image stitching. The long-standing interest in this problem has led to various approaches. Early methods like SIFT[22] and ORB[23] laid the groundwork with handcrafted keypoint descriptors. These algorithms created robust descriptors invariant to scale, rotation, and partial occlusion, making them effective for many tasks despite sensitivity to image illumination changes. Recently, deep learning techniques revolutionized keypoint matching by learning discriminative features from large datasets. CNNs, like the SuperPoint model[24], have been widely adopted for keypoint detection and matching, learning detection and descriptor generation end-to-end. Deep learning approaches offer improved robustness and generalization over handcrafted methods. Recent advances use attention mechanisms to improve keypoint matching in challenging environments. Networks like SuperGlue[25] use graph neural networks with self- and cross-attention to establish robust keypoint correspondences. Transformers in keypoint matching improve context understanding, handling complex matching scenarios more proficiently.
Inspired by SuperPoint and Superglue, we consider each object in the scene as a keypoint, and compress the keypoint features into a single feature vector using the proposed dynamic channel graph attention (DCGA) mechanism. Then, we establish object-to-object matching relationships using the proposed OMM.
@@ II-C Unsupervised Anomaly Detection
Current unsupervised anomaly detection approaches are typically classified into two categories: methods targeting structural anomalies and those targeting logical anomalies. Most current anomaly detection methods target structural anomalies, using robust feature extraction networks to obtain high-level semantic features of a test image. The distance between features of test images and anomaly-free images is calculated, identifying areas with large distances as anomalies. Existing methods typically use deep CNN models like ResNet[26] and EfficientNet[27] for feature extraction. For example, Roth et al.[10] proposed Patchcore to detect anomalies on objects with a concise background. They selected ResNet-50[26] pretrained on ImageNet[28] as the feature extractor. However, these methods struggle with logical anomalies since the local structures within the image appear normal. Therefore, more researchers are focusing on logical anomaly detection. For instance, Zhang et al. proposed DSKD[15], and Batzner introduced EfficientAD[29]. Both methods use the Student-Teacher network, detecting anomalies by comparing teacher and student differences. Using a local-global branching approach, Yao et al. presented GLCF[13], while Bergmann et al. proposed GCAD[11] to understand the logical semantic constraints of the entire image scene. Guo et al. presented THFR[14], based on template-guided reconstruction. These methods have shown promise in detecting certain logical anomalies, such as missing components and unexpected excess. However, their performance drops significantly with more complex logical anomalies, such as misordering, mismatches, and haphazard object arrangements.
The proposed SAM-LAD builds a zero-shot method without strenuous efforts to comprehend the logical semantics of the entire scene. By introducing SAM and proposing OMM and AMM, a global logical anomaly detection system was developed, effectively addressing the limitations of existing methods mentioned above.
@@ III-E1 Dynamic Channel Graph Attention
This graph structure consists of two parts. The first part targets each object’s feature vector $(1,C)$. Specifically, two independent $C\times C$ matrices, that is, $A_{0}$ and $A_{1}$, constitute the adjacency matrices, representing the dependency relationships among channel vertices. $A_{0}$ is generated by a 1-D convolution layer and a Softmax function. It is a self-attention-based diagonal matrix designed to suppress irrelevant features. $A_{1}$ is a predefined identity matrix, representing only the vertex itself, and requires normalization. Thus, the adjacency matrix can be represented as $A=A_{0}\times A_{1}$.
@@
$A_{2}$ $=\begin{bmatrix}0&S_{12}&\cdots&S_{1j}\\ S_{21}&0&\cdots&S_{2j}\\ \vdots&\vdots&\ddots&\vdots\\ S_{i1}&S_{i2}&\cdots\ &S_{ij}\\ \end{bmatrix}.$ (4)
Where $S_{ij}^{\prime}$ represents the similarity between two object vertex in the M and $S_{ij}$ is its normalization, denoted as:
$S_{ij}=\frac{S_{ij}^{\prime}}{\sum_{k=1}^{M}S_{ik}},S_{ij}^{\prime}=\hat{v}_{i}\cdot\hat{v}_{j}=\sum_{d=1}^{C}\hat{v}_{id}\cdot\hat{v}_{jd}.$ (5)
Note that, when generating the adjacency matrix, the weights updating is based solely on the similarity between one object vector and others, self-similarity is set to zero, that is, the diagonal of $A_{2}$ is zero.
@@ III-F Anomaly Measurement Module
Specifically, based on the indices from the $\mathbf{P}$ and trash bins, we calculate the difference in feature distribution for the object feature maps that were successfully matched with the corresponding maps in the $k$ reference images to create the matching score map. Similarly, we calculate the difference in feature distribution for the unmatched object feature maps with the most closely matched object feature maps in the $k$ reference images to create the non-matching score map. The final score map is obtained by adding these two score maps together. This approach allows for both the detection of logical anomalies and detailed checks for structural anomalies within each object.
@@ IV-B Evaluation metrics
IV-C Implementation Details
IV-D Comparison With the State-of-the-Art Models
@@ IV-D2 MVTec LOCO AD
The comparative results are presented in Table II. Compared to the best baseline method and the SoTA method, our framework improves +25.6 and +0.7 on image-level AUROC, +45.0 and +5.4 on pixel-level sPRO, respectively.
Note that, among all the comparison methods, only our approach implemented zero-shot while simultaneously maintaining the best detection performance. Furthermore, Fig. 8 displays several detection results of SAM-LAD on both structural and logical anomalies within the MVTec LOCO AD dataset, illustrating that our framework is adept at precisely pinpointing both types of anomalies.
@@
IV-D3 MVTec AD
Beyond affirming the performance of our SAM-LAD in detecting logical anomalies, we further appraise its capability to identify structural anomalies on the MVTec AD dataset, which exclusively encompasses structural anomalies. To be specific, given that each subset within the MVTec AD dataset features a scene with a single object against a regular background, we facilitate the segmentation into this single object by employing the SAM filter for subsequent matching and computation. Owing to the one-to-one fixed matching paradigm, the ultimate detection results are directly yielded by the AMM, thereby serving as an indirect assessment of the efficacy of our proposed AMM.
IV-D4 DigitAnatomy
IV-E Further Analysis
The outstanding logical anomaly detection results of our framework are based on the good performance of the Object Matching Model (OMM). Therefore, to verify the reasons behind the exceptional performance, for each category in MVTec LOCO AD and DigitAnatomy dataset, we compute the precision, recall, and F1 score of OMM’s matching results, which are shown in Table V.
The results demonstrate the OMM’s strong matching capability and validate its effectiveness. However, the matching accuracy for the pushpins and splicing connectors categories drops significantly. This occurs because each object is very similar in these scenarios, posing a challenge for the OMM. Instead of using unmatched object masks as the anomaly score, we designed the AMM (Anomaly Measurement Module) to further detect anomalies. This design reduces the SAM-LAD’s dependency on matching accuracy, as shown by the final anomaly detection results in Section IV-D.
TABLE V: The Performance of the OMM’s Matching capability in MVTec LOCO AD dataset and Digitanatomy Dataset.
[…]
MVTec LOCO AD
Breakfast box 99.2% 98.8% 99.0%
Juice Bottle 99.5% 98.3% 98.9%
Pushpins 89.0% 86.9% 87.9%
Screw bag 99.5% 98.9% 99.2%
Splicing connectors 91.1% 87.9% 89.5%
Mean 95.7% 94.2% 96.5%
Digitanatomy Dataset
Mean 96.5% 95.1% 95.8%
@@ IV-F1 Impact of the Backbone
We use the pre-trained backbone as the feature extractor in the proposed SAM-LAD. In this study, we primarily considered three different pre-trained backbone networks: the Wide-ResNet50 of the CNN type, the Swin-transformer of the ViT type, and the DINOv2 of the ViT type. The main results are presented in Table VI, where it is distinctly that DINOv2 manifests the most superior performance. Therefore, we adopt DINOv2 as the feature extractor for the proposed framework.
TABLE VI: Ablation Experiments of the Different Backbone on the MVTec LOCO AD Dataset.
[…]
@@
IV-F2 Impact of the FeatUp Configuration
After obtaining the feature maps from the images, we use the FeatUp operation to upsample these maps and restore the lost spatial information. In the study of FeatUp[18], the authors proposed five upsampling configurations: 2$\times$, 4$\times$, 8$\times$, 16$\times$, and 32$\times$, shown in Fig. 13. To verify FeatUp’s efficacy and assess the impact of different upsampling configurations on anomaly detection performance, we conduct an ablation study on the upsampling factors.
@@ IV-F2 Impact of the FeatUp Configuration
The main results are shown in Table VII. Compared to the original feature maps, FeatUp significantly enhances detection capabilities. Specifically, as the upsampling factor increases, SAM-LAD’s detection and segmentation performance improves. However, upsampling the feature maps by 16$\times$ and 32$\times$ greatly increases the model’s FLOPs and drastically reduces inference speed, which is impractical for industrial use. Therefore, to balance detection performance with real-time inference needs, we chose an 8$\times$ upsampler to restore lost spatial information in the feature maps for subsequent computations.
TABLE VII: Ablation experiments of the Different FeatUp’s Upsampling Factors on the MVTec LOCO AD Dataset.
@@
TABLE VIII: Ablation Experiments of the Different Feature Compress Methods on the MVTec LOCO AD Dataset.
@@
IV-F3 Impact of the DCGA
To validate the effectiveness of DCGA, we conducted an ablation experiment. DCGA was designed to extract high-dimensional features into a single vector, simplifying the similarity computation between objects. Pooling is the most widely used method for feature extraction. We compared DCGA with global average pooling (GAP) and global max pooling (GMP), as shown in Table VIII. GMP outperforms GAP because only certain areas of the object feature map have discernible values, while the background is mostly null. Therefore, GMP is more effective in capturing object features in such scenarios. However, simply selecting the maximum value for pooling is insufficient. The introduction of DCGA greatly enriched the description of object features. Using graph neural network principles, DCGA extracts effective object features from high-dimensional channels and captures feature interrelationships among objects in the same scene. This substantially improves the matching capabilities of the OMM, confirming the validity of DCGA.